Translate each of the following word expressions into an Algebraic Expression
1.Twelve more than five times a number
2.Six times the sum of a number and four
3.Eight subtracted from two times a number.
4.The quotient of one less than a number and twice a number
6.The product of a number and four increased by the number
5.Eleven times the difference of a number and three is equal to twice the number
7.If seven times a number is reduced by nine,the result is ten less than the number
Thank You
Answer #1
1) 5x+12 2) 6(x+4) 3) 2x-8 4) 2x/x-1 5) 11(x-3)=2x 6) 4x+x 7) 7x-9=x-10
answered by: daryl